The bulk modulus of water is 2.2 x 10^9 N/m2 . Find the pressure necessary to change the volume of water by 1.0 %. [1 bar = 105 N/m2] 
 
(a) 2.2 bar 
(b) 110 bar 
(c) 220 bar 
(d) 220 mbar 

Grade:12

1 Answers

Somi
6 years ago
K(bulk modulus)K= -P÷fractional change in volumeK=2.2×10^9=-P÷1% =-P÷.01 :. P= 2.2×10^7Or P= 220×10^5I.e. 220 bar